A School has 575 students. 44% are boys. How many girls are in the school
1
Expert's answer
2012-12-11T09:27:30-0500
There are

B = 575 * (44%/100%) = 575 * 0.44 = 253

boys at the school. Therefore, the number of girls is

G = 575 - 253 = 322.